How many moles of nitrogen, N

N, are in 81.0 g

g of nitrous oxide, N

2

O

N2O?


1
Expert's answer
2021-02-16T07:43:16-0500

Solution:

"mole=\dfrac{m}{Mr}=\dfrac{81\;g}{44(\frac{g}{mole})}=1.84\;moles\;of\;N_2O"


Since, clearly, there are 2 moles of nitrogen atoms per mole of nitrous oxide, we simply double the molar quantity of nitrous oxide:


1.84 moles * 2 = 3.68 moles of Nitrogen.